OpenStudy (anonymous):

Solve: 40x2 - 77x = -36 A. {6/5, 3/4} B. {-6/5, -3/4} C. {4/5, 9/8} D. {-4/5, -9/8}

OpenStudy (anonymous):

start with \[40 x^2-77 x+36 = 0\] and by some complete miracle this actually factors

OpenStudy (anonymous):

not sure how you are supposed to get it, but it is \\[(5 x-4) (8 x-9)=0\] and now the zeros are easy to find

OpenStudy (anonymous):

x=4/5,9/8 so it is c